OpenStudy (anonymous):

What is the circumference of the circle whose equation is (x - 9)2 + (y - 3)2 = 64? 8p 64p 16p 4p

OpenStudy (stamp):

Circumference is defined as C = 2pi r The missing piece is knowing the radius. Can you determine the radius of our circle in question? Perhaps you can find the r value from the circle equation. Once you know the radius, you can substitute the radius r into your C = 2pi r and determine your circumference

OpenStudy (anonymous):

i have no clue on how to find the radius

OpenStudy (stamp):

This is the general form of a circle equation\[(x-a)^2 + (y-b)^2 = r^2\]

OpenStudy (anonymous):

ohh so its 8p?

OpenStudy (stamp):

\[r^2=64\]so\[r=8\] If C = 2pi r, and r = 8 What is C?

OpenStudy (anonymous):

16?